OpenStudy (anonymous):

5. What is one value of “k” for which the homogeneous system has infinitely many solutions? System: (5-k)x+y=0 2x+ (4-k)=0

OpenStudy (freckles):

or those really the equations?

OpenStudy (freckles):

are you missing a y in the second one?

OpenStudy (freckles):

I think you mean (5-k)x+y=0 2x+(4-k)y=0

OpenStudy (freckles):

in that case find the slopes of both lines and set them equal --- if we have same slope and same y-intercept then we have infinitely many solutions

OpenStudy (anonymous):

a1/a2=b1/b2=c1/c2 use this

OpenStudy (anonymous):

Im doing this with matrices, I know how to do it with equations, and I know I need to get a row of "0s" to get infinitely many solutions, and I also know what it looks like on a graph. I need to know how to do it for my Linear Algebra class, and I can't seem to be able to use elementary row operations to find k.

OpenStudy (freckles):